# Hausdorff Dimension of Union of Lines Covering a Curve: Applications to Mathematical Physics

### 26 Mar 2026 v2

@@ Abstract
We prove that for any nonlinear $f\in C^{1,\alpha}([0,1])$, the union of lines covering its graph over a sufficiently large full measure subset has a Hausdorff dimension of at least $1+\alpha$, and this dimension bound is sharp. We then apply these geometric results to mathematical physics, proving that light rays forming a differentiable caustic in the plane must illuminate a 2D region, and that spacetime observability sets for conservation laws with $\alpha$-Hölder initial wave speeds possess a dimension of at least $\alpha$. Finally, we establish a measure theoretic result: For a continuous differentiable function $f$ of which derivative is non-constant of bounded variation, if the union of some family of lines that cover $\operatorname{graph}(f)$ has Hausdorff dimension less than $2$, then the distributional derivative $f^{\prime\prime}$ is a singular measure.
@@ 1 Introduction and Background
Throughout this article, we denote the Hausdorff dimension of a subset $E$ of the Euclidean space $\mathbb{R}^{d}$ by $\dim_{\mathcal{H}}(E)$.
@@
Figure 1.1: The Sierpiński triangle of Hausdorff dimension $\alpha=\log(3)/\log(2)\approx 1.585$.
When the covering family is permitted to include tangent lines, a striking dichotomy emerges based on the differentiability of the curve. Cumberbatch et al. [1] recently demonstrated that if a nonlinear curve is twice-differentiable, the union of its covering lines must still possess a Hausdorff dimension of exactly 2. Astonishingly, they proved that this rigidity breaks down at lower regularities: it is possible to construct a continuously differentiable strictly convex curve that can be perfectly covered by a family of lines of which union has dimension $1$.
We make five main contributions to this line of inquiry. First, in Section 2, we significantly strengthen a related result by Cumberbatch et al. regarding the tangent lines of differentiable functions. We prove that the union of tangent lines of a nonlinear differentiable function does not merely possess a Hausdorff dimension of 2, but must in fact contain an open disk.
@@
Fourth, in Section 5, we translate our purely geometric results into the realm of nonlinear partial differential equations, providing a novel dimension bound for the observability of characteristic lines in general first-order scalar conservation laws.
Finally, in Section 6, we focus on the measure theory of functions of bounded variation. Utilizing the Lebesgue decomposition theorem and the area formula, we prove that for a nonlinear $C^{1}$ function $f$ of which derivative is of bounded variation, if the union of some family of lines that cover $\operatorname{graph}(f)$ has Hausdorff dimension less than $2$, then the distributional derivative $f^{\prime\prime}$ is a singular measure.
@@ 2 Graphs of Differentiable Functions
In this section, we show that this geometric structure is far more rigid. We establish that the union of these tangent lines cannot be a highly porous or purely fractal set of dimension 2; rather, the nonlinearity of the curve forces the tangent lines to sweep out positive area, guaranteeing that their union contains an open disk, or equivalently, a set with non-empty interior. We first establish a crucial connectivity property.
@@ 4 Hölder and Sobolev Spaces
Notice that, the class of analytic sets is closed under countable unions and intersections, also under continuous images or inverse images. Another core property of analytic sets is that they are always Lebesgue measurable.
For our purpose, the most important hallmark of analytic sets is revealed by the following classical theorem.
@@ 5 Applications to Conservation Laws
It is physically tempting to look for a trivial observability set $\Lambda$ of dimension zero. Indeed, if the initial wave speed function $g(x)=F^{\prime}(f(x))$ is purely linear, the characteristic lines can perfectly focus into a single spacetime focal point, creating an instantaneous shock. Placing a single sensor at this exact focal point intersects every characteristic simultaneously, implying that full observability can be achieved with a discrete set of dimension 0.
@@
However, our corollary rigorously proves that this perfect 0-dimensional focusing is entirely an artifact of a linear initial wave speed. When the physical interaction between the initial data $f(x)$ and the flux function $F(u)$ produces a wave speed $g(x)$ that is nonlinear and parameterized by an intermediate Hölder regularity $C^{1,\alpha}$, the varying slopes of the characteristics physically prevent them from perfectly aligning into a single focal point. Instead, the resulting shocks and caustics are necessarily smeared across spacetime. Consequently, any sensor network or control domain $\Lambda$ capable of intersecting all information pathways must maintain a minimal Hausdorff dimension of $\alpha$. This reveals a deep and necessary connection between the fractal dimension of spacetime observatory and the regularity of the physical waves.
@@ Appendix A Methods from Topology
We begin by establishing a strengthened variation of the open mapping theorem. Specifically, we prove that a continuous map possessing a full rank differential at merely a single point is topologically forced to map onto a set with a non-empty interior.
@@
Then, $\Psi_{\mathbf{v}}$ is continuous, and $\Psi_{\mathbf{v}}(\mathbf{u})=\mathbf{u}$ if and only if $J^{-1}(f(\mathbf{u})-\mathbf{v})=0$, that is, $\mathbf{v}=f(\mathbf{u})$. Moreover we have
@@
Take any $\mathbf{u}\in A$. Then $\mathbf{v}=f(\mathbf{u})$ and hence $\Psi_{\mathbf{v}}(\mathbf{u})=\mathbf{u}$. Therefore by the estimation (*A), we immediately have
@@ Appendix B Generalization to Hypersurfaces
$\dim_{\mathcal{H}}(\{\mathbf{y}\in\mathbb{R}^{m}:(\mathbf{x},\mathbf{y})\in E\})\geq\beta$
for all $\mathbf{x}\in A$, then $\dim_{\mathcal{H}}(E)\geq\alpha+\beta$.
@@ Acknowledgement
The author is deeply grateful to Tamás Keleti for drawing the author’s attention to this fascinating topic and for inspiring this line of inquiry.
Furthermore, the author is also grateful to Tamás Keleti for pointing out a mathematical inaccuracy occurred in section 3, and for giving an idea on how to fix that.

### 31 Mar 2026 v3

@@ Abstract
We prove that for any nonlinear $f\in C^{1,\alpha}([0,1])$, the union of lines covering its graph has a Hausdorff dimension of at least $1+\alpha$, and this dimension bound is sharp. We then apply these geometric results to mathematical physics, proving that spacetime observability sets for conservation laws with $\alpha$-Hölder initial wave speeds possess a dimension of at least $\alpha$. Finally, we prove that if an absolutely integrable vector field $v$ on the boundary of a polyhedron exhibits a strictly positive total flux, then the union of the line field spanned by $v$ possesses a Hausdorff dimension of 3.
@@ 1 Introduction and Background
Figure 1.1: The Sierpiński triangle of Hausdorff dimension $\alpha=\log(3)/\log(2)\approx 1.585<2$. Throughout this article, we denote the Hausdorff dimension of a subset $E$ of the Euclidean space $\mathbb{R}^{d}$ by $\dim_{\mathcal{H}}(E)$.
When the covering family is permitted to include tangent lines, a striking dichotomy emerges based on the differentiability of the curve. Cumberbatch et al. recently demonstrated in [1] that if a nonlinear curve is twice-differentiable, the union of its covering lines must still possess a Hausdorff dimension of exactly 2. Astonishingly, they proved that this rigidity breaks down at lower regularities: it is possible to construct a continuously differentiable strictly convex curve that can be perfectly covered by a family of lines of which union has dimension $1$.
We make six main contributions to this line of inquiry. First, in Section 2, we significantly strengthen a related result by Cumberbatch et al. regarding the tangent lines of differentiable functions. We prove that the union of tangent lines of a nonlinear differentiable function does not merely possess a Hausdorff dimension of 2, but must in fact contain an open disk.
@@
Fourth, in Section 5, we translate our geometric results into the realm of nonlinear partial differential equations, providing a novel dimension bound for the observability of characteristic lines in general first-order scalar conservation laws.
Fifth, in Section 6, we focus on the measure theory of functions of bounded variation. Utilizing the Lebesgue decomposition theorem and the area formula, we prove that for a nonlinear $C^{1}$ function $f$ of which derivative is of bounded variation, if the union of some family of lines that cover $\operatorname{graph}(f)$ has Hausdorff dimension less than $2$, then the distributional derivative $f^{\prime\prime}$ is a singular measure.
Finally, in Section 7, as a corollary of Wang and Zahl’s recent work in Kakeya problem, we prove that any absolutely integrable vector field defined on the boundary of a polyhedron that yields a strictly positive total flux forces the corresponding line field to sweep out a set of full Hausdorff dimension.
@@ 2 Graphs of Differentiable Functions
In this section, we show that this geometric structure is far more rigid. We establish that the union of these tangent lines cannot be a highly porous or fractal set of dimension 2; rather, the nonlinearity of the curve forces the tangent lines to sweep out positive area, guaranteeing that their union contains an open disk, or equivalently, a set with non-empty interior. We first establish a crucial connectivity property.
@@ 4 Hölder and Sobolev Spaces
Notice that the class of analytic sets is closed under countable unions and intersections, as well as under continuous images or inverse images. Another core property of analytic sets is that they are always Lebesgue measurable.
For our purposes, the most important hallmark of analytic sets is revealed by the following classical theorem.
@@ 5 Applications to Conservation Laws
It is physically tempting to look for a trivial observability set $\Lambda$ of dimension zero. Indeed, if the initial wave speed function $g(x)=F^{\prime}(f(x))$ is linear, the characteristic lines can perfectly focus into a single spacetime focal point, creating an instantaneous shock. Placing a single sensor at this exact focal point intersects every characteristic simultaneously, implying that full observability can be achieved with a discrete set of dimension 0.
@@
However, our corollary rigorously proves that this perfect 0-dimensional focusing is entirely an artifact of a linear initial wave speed. When the physical interaction between the initial data $f(x)$ and the flux function $F(u)$ produces a wave speed $g(x)$ that is nonlinear and parameterized by an intermediate Hölder regularity $C^{1,\alpha}$, the varying slopes of the characteristics physically prevent them from perfectly aligning into a single focal point. Instead, the resulting shocks and caustics are necessarily smeared across spacetime. Consequently, any sensor network or control domain $\Lambda$ capable of intersecting all information pathways must maintain a minimal Hausdorff dimension of $\alpha$. This reveals a deep and necessary connection between the fractal dimension of spacetime observatories and the regularity of the physical waves.
@@ 6 Functions of Bounded Variation
7 Main Theorem on Polyhedrons
We have already explored the dimensional rigidity of line families covering planar curves. One may ask how these geometric constraints manifest in higher-dimensional ambient spaces, particularly in $\mathbb{R}^{3}$.
The study of unions of lines in $\mathbb{R}^{3}$ is inextricably linked to the notorious Kakeya set conjecture, which asserts that any compact set containing a unit line segment in every direction must possess a Hausdorff and Minkowski dimension of exactly 3. While the two-dimensional case was resolved by Davies in his seminal work [3], the three-dimensional case stood as one of the most prominent open problems in geometric measure theory for decades. See [6] and [10] for an introduction to the Kakeya conjecture and a survey of historical progress on the problem.
Recently, in a monumental breakthrough, Wang and Zahl definitively resolved the Kakeya set conjecture in three dimensions. By utilizing a refined multi-scale decomposition and proving that unions of tubes satisfying specific conditions must attain almost maximal volume, they closed the dimensional gap completely. For our purposes, their result provides the ultimate geometric rigidity required for line families whose directions do not collapse into a null set. We state this continuous formulation of their theorem as follows:
Theorem 7.1 (Wang-Zahl, 2025).
Let $\Omega$ be a measurable subset of the unit $2$-sphere
$\mathbb{S}^{2}:=\{\mathbf{x}\in\mathbb{R}^{3}:\|\mathbf{x}\|=1\},$
and $F\colon\Omega\rightarrow\mathbb{R}^{3}$ a measurable vector field. If $\Omega$ enjoys positive Haar measure, then the set
$\bigcup_{\mathbf{x}\in\Omega}\{F(\mathbf{x})+\mathbf{x}t:t\in[0,1]\}$
has Hausdorff dimension $3$.
With this profound dimensional bound established for families of lines possessing a positive measure of directions, we shift our focus to vector fields defined on the two-dimensional boundaries of three-dimensional spatial domains. Specifically, we examine the case where the domain is a polyhedron. The geometric flatness of polyhedral faces provides a rigid algebraic structure that interacts remarkably well with fluid flux. If an absolutely integrable vector field exhibits a strictly positive total flux across the polyhedral boundary, it must inherently point transversally outward on a subset of a flat face with strictly positive two-dimensional Lebesgue measure.
By exploiting projective duality, we can map this spatial plane to the plane at infinity. This projective linear transformation elegantly converts the positive measure of spatial base points directly into a positive measure of directions, thereby seamlessly satisfying the exact hypotheses of the Wang-Zahl theorem on Kakeya sets. This enables us to establish our main geometric constraint for polyhedral boundaries:
Theorem 7.2.
Let $P$ be a polyhedron in $\mathbb{R}^{3}$, and $v\in L^{1}(\partial P)^{\oplus 3}$ an absolutely integrable vector field. Suppose that the total flux of $v$ through $\partial P$ is positive. Then the set
$\bigcup_{\mathbf{x}\in\partial P}\{\mathbf{x}+v(\mathbf{x})t:t\geq 0\}$
has Hausdorff dimension $3$.
Proof.
By Dirichlet’s Schubfachprinzip and the assumption on the total flux, it suffices to prove that, for a measurable subset $\Omega$ of $\mathbb{R}^{2}$ and measurable functions $f,g\colon\Omega\rightarrow\mathbb{R}$, the set
$E:=\bigcup_{(x,y)\in\Omega}\{(x+f(x,y)t,y+g(x,y)t,t):t>0\}$
has Hausdorff dimension $3$ whenever $\Omega$ is not a null set.
Consider the projective linear transformation
$\varphi\colon\{(x,y,z)\in\mathbb{R}^{3}:z\neq 0\}\rightarrow\\ \{(x,y,z)\in\mathbb{R}^{3}:z\neq 0\}$
defined by
$\varphi(x,y,z):=\left(\frac{x}{z},\frac{y}{z},\frac{1}{z}\right).$
Take any $(x,y)\in\Omega$. Applying $\varphi$ to the ray
$L:=\{(x+f(x,y)t,y+g(x,y)t,t):t>0\}$
yields that
$\varphi(L)=$ $\{(f(x,y)+x/t,g(x,y)+y/t,1/t):t>0\}$
$=$ $\{(f(x,y)+xt/\rho,g(x,y)+yt/\rho,t/\rho):t>0\}$
for $\rho(x,y):=\sqrt{x^{2}+y^{2}+1}$.
Since $\{(x,y,1)/\rho(x,y):(x,y)\in\Omega\}$ is a measurable set of positive measure, as $\Omega$ is, the set $\varphi(E)$ has Hausdorff dimension $3$ by Theorem 7.1. Finally, because $\varphi$ is a smooth involution, we conclude that $\dim_{\mathcal{H}}(E)=3$. ∎
@@ Appendix A Methods from Topology
We begin by establishing a strengthened variation of the open mapping theorem. Specifically, we prove that a continuous map possessing a full rank differential at merely a single point is forced to map onto a set with a non-empty interior.
@@
Then, $\Psi_{\mathbf{v}}$ is continuous, and $\Psi_{\mathbf{v}}(\mathbf{u})=\mathbf{u}$ if and only if $J^{-1}(f(\mathbf{u})-\mathbf{v})=0$, that is, $\mathbf{v}=f(\mathbf{u})$. Moreover, we have
@@
Take any $\mathbf{u}\in A$. Then $\mathbf{v}=f(\mathbf{u})$ and hence $\Psi_{\mathbf{v}}(\mathbf{u})=\mathbf{u}$. Therefore by the estimate (*A), we immediately have
@@ Appendix B Generalization to Hypersurfaces
$\dim_{\mathcal{H}}(\{\mathbf{y}\in\mathbb{R}^{m}:(\mathbf{a},\mathbf{y})\in E\})\geq\beta$
for all $\mathbf{a}\in A$, then $\dim_{\mathcal{H}}(E)\geq\alpha+\beta$.
@@ Acknowledgement
The author is deeply grateful to Tamás Keleti for drawing the author’s attention to this fascinating topic and for inspiring this line of inquiry. Also, the author would like to express sincere gratitude to Weiyi Zhang for very useful and inspiring discussions.
Furthermore, the author is also grateful to Tamás Keleti for pointing out a mathematical inaccuracy that occurred in Section 4, and for suggesting an idea on how to fix that.